Competency Question 5. Calculate the molar solubility of salt in pure water. The Ksp value is The instructors will choose a salt such as CaCO3, Pbl2, Hg2Cl2, Ag3PO4, or Ca3(PO4)2. The instructors will give a numerical value for Ksp appropriate to the salt. Your full credit answer must show (1) the balanced equilibrium reaction equation, (2) the Ksp equilibrium constant expression written in terms of x, (3) essential mathematical work, and (4) the molar solubility as the final answer. You can use an ICE table or the shortcut method to derive the Ksp expression. For practice, use the five listed salt formulas and assume a Ksp value of 1.6x10 8
Answer #1

Given Ksp for sample = 1.6 * 10-8

(1)

CaCO3 (s) \rightleftharpoons Ca2+ (aq.) + CO32- (aq.)

If solubility of CaCO3 is S mol/L, then [Ca2+] = [CO32-] = S mol/L

Ksp = [Ca2+][CO32-]

1.6 * 10-8 = (S)(S)

S = solubility of CaCO3 = 1.26 * 10-4 M

(2)

PbI2 (s) \rightleftharpoons Pb2+ (aq.) + 2 I- (aq.)

If Solublity of PbI2 = S mol/L, then, [Pb2+] = S mol/L, [I-] = 2S mol/L

Ksp = [Pb2+][I-]2

1.6 * 10-8 = (S)(2S)2

S = 1.59 * 10-3 M

(3)

Hg2Cl2 (s) \rightleftharpoons Hg22+ (aq.) + 2 Cl- (aq.)

If Solublity of Hg2Cl2 = S mol/L, then, [Hg22+] = S mol/L, [Cl-] = 2S mol/L

Ksp = [Hg22+][Cl-]2

1.6 * 10-8 = (S)(2S)2

S = 1.59 * 10-3 M

(4)

Ag3PO4 (s) \rightleftharpoons 3 Ag+ (aq.) + PO43- (aq.)

If solubility of Ag3PO4 is S mol/L, then, [Ag+] = 3S mol/L and [PO43-] = S mol/L

Ksp = [Ag+]3[PO43-]

1.6 * 10-8 = (3S)3(S)

S = 4.93 * 10-3 M

(5)

Ca3(PO4)2 (s) \rightleftharpoons 3 Ca2+ (aq.) + 2 PO43- (aq.)

If the solubility = S mol/L, then, [Ca2+] = 3S mol/L and [PO43-] = 2S mol/L

Ksp = [Ca2+]3[PO43-]2

1.6 * 10-8 = (3S)3(2S)2

S = 1.71 * 10-2 M